# Game Theory in Cosmology

### v1

@@ 4 Cosmological Teleodynamics Framework
$\mathcal{P}[\Gamma]\propto\exp[-\beta\mathcal{A}[\Gamma]-\alpha\mathcal{K}[\Gamma]]$ (20)
@@
$\mathcal{A}[\Gamma]=\int dt\left[\frac{p^{2}}{2a^{2}m}+m\Psi(x,t)\right]$ (21)
@@
$\mathcal{K}[\Gamma]=\int dt\,\Phi\left(x(t),p(t);f,\mathcal{E}\right)$ (22)
@@
$\dot{x}=\frac{p}{a^{2}m}$ (23)
$\dot{p}=-\nabla\Psi(x,t)-\alpha\nabla\Phi(x,p,t)-Hp$ (24)
@@
This is helpful as it shows that the dynamics experiences an additional force proportional to the gradient of the bias functional. This force is not sourced from new matter fields but rather represents a coarse-grained structural response of the system to its own past evolution and environment. As a result, the presence of $\alpha\nabla\Phi$ alters the gravitational streaming of matter, modifies the effective potential, and influences the background expansion.
By promoting the trajectory equations to a kinetic description, one obtains the teleodynamic Boltzmann equation, from which the modified continuity, Euler, and growth equations follow. One can decompose the bias functional into its background part and fluctuations,
$\Phi(t,x)=\bar{\Phi}(t)+\varphi(t,x)$ (25)
where $\bar{\Phi}(t)$ contributes a homogeneous energy density to the Friedmann equations and $\varphi(t,x)$ modifies the Poisson equation through an effective clustering density. This decomposition is central, as it provides a view of the dark sector and, therefore, the bias functional and the maximum-caliber weight provide the unifying mathematical structure underlying all teleodynamic effects in cosmology.
@@ 4.1 Boltzmann Dynamics and Galactic Phase-space Evolution
$\frac{\partial f}{\partial t}+\frac{\mathbf{p}}{a^{2}m}\cdot\nabla_{\mathbf{x}}f-\nabla_{\mathbf{x}}\Psi\cdot\nabla_{\mathbf{p}}f-H\,\mathbf{p}\cdot\nabla_{\mathbf{p}}f=0$ (26)
@@
$\Phi(t,\mathbf{x})=\bar{\Phi}(t)+\varphi(t,\mathbf{x})$ (25)
@@
$\mathbf{a}=-\nabla\Psi-\alpha\,\nabla\Phi$ (27)
@@
$\frac{\partial f}{\partial t}+\frac{\mathbf{p}}{a^{2}m}\cdot\nabla_{\mathbf{x}}f-\left(\nabla_{\mathbf{x}}\Psi+\alpha\nabla_{\mathbf{x}}\Phi\right)\cdot\nabla_{\mathbf{p}}f-H\,\mathbf{p}\cdot\nabla_{\mathbf{p}}f=0$ (28)
@@
$\dot{\delta}+\frac{1}{a}\nabla\cdot\left[(1+\delta)u\right]=0$ (29)
@@
$\dot{u}+Hu+\frac{1}{a}(u\cdot\nabla)u=-\frac{1}{a}\nabla\Psi-\frac{1}{a(1+\delta)}\nabla\cdot P-\frac{\alpha}{a}\nabla\Phi$ (30)
@@
$\dot{\delta}(k,t)=-\frac{1}{a}\,\theta(k,t),$ (31)
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga\bar{\rho}\,\delta(k,t)=-\frac{c_{s}^{2}k^{2}}{a}\,\delta(k,t)-\frac{\alpha k^{2}}{a}\,\Phi_{k}(t)$ (32)
@@ 4.2 Friedmann Dynamics and Dark Energy
$\dot{\rho}+3H(\rho+p)=0$ (36)
these relations fully determine the FLRW background given a set of matter components with equations of state. One immediate consequence of (34)-(35) is that cosmic acceleration
$\frac{\ddot{a}}{a}>0,$ (37)
@@
$\rho_{\text{TD}}(t)=\alpha\bar{\Phi}(t),\quad p_{\text{TD}}(t)=\alpha\Pi_{\text{TD}}(t)$ (40)
Note that no additional fields are required for this term to appear. It follows directly from the statistical and systemic coarse-graining of matter in a gravitationally correlated environment and represents the macroscopic imprint of dynamical memory and nonlocal structure. One can derive (38) and (39) directly by inserting the modified stress tensor
@@
$S_{\mathrm{eff}}[g_{\mu\nu}]=\int d^{4}x\,\sqrt{-g}\left[\mathcal{L}_{\mathrm{matter}}-\alpha\,\Phi\right]$ (41)
@@
$\frac{\ddot{a}}{a}=-\frac{1}{6M_{P}^{2}}[\rho+3p+\rho_{\text{TD}}+3p_{\text{TD}}]>0$ (42)
@@
$w_{\text{TD}}\equiv\frac{p_{\text{TD}}}{\rho_{\text{TD}}}=\frac{\Pi_{\text{TD}}(t)}{\bar{\Phi}(t)}\approx-1$ (43)
@@
Instead of introducing new scalar fields, potentials, or hidden matter sectors in our theory [52, 53, 54], one simply recognizes that the universe is not a system that relaxes to a gas-like equilibrium or one whose expansion is governed solely by local microphysics. The long-range nature of gravity and the persistent structures generated by cosmic evolution imply that the background expansion should include contributions that have nonlocal correlations and path dependence. The teleodynamic energy density $\rho_{\text{TD}}$ arises precisely from these considerations and provides an emergent explanation for the cosmic acceleration. The term $\rho_{\mathrm{TD}}=\alpha\bar{\Phi}$ arises not from a new microscopic field but from the coarse-grained effect of self-gravitating correlations and memory, which accumulate as the Universe forms persistent structures and departs more strongly from ergodicity. In this sense, we see that late-time acceleration emerges naturally as the Universe becomes increasingly non-thermal and the expansion is governed by an effective potential generated by the ensemble of long-lived gravitational correlations. Teleodynamics in this sense provides a statistical and systemic explanation for cosmic acceleration grounded in the physics of memory-bearing, nonlocally correlated matter rather than in additional dark sectors or fine-tuned vacuum energies.
@@ 4.3 Dark Matter and Structure Growth
$\Psi_{\text{eff}}(t,x)=\Psi(t,x)+\alpha\varphi(t,x)$ (44)
@@
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}\bar{\rho}\delta+\alpha\nabla^{2}\varphi$ (46)
@@
$\rho_{\text{TD}}(t,x)=\frac{\alpha}{4\pi Ga^{2}}\nabla^{2}\varphi(t,x)$ (47)
@@
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}[\bar{\rho}\delta+\rho_{\text{TD}}]$ (48)
@@
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga\bar{\rho}\delta(k,t)=-\frac{\alpha k^{2}}{a}\varphi_{k}(t)$ (49)
@@
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga[\bar{\rho}\delta(k,t)+\rho_{\text{TD}}(k,t)]=0$ (50)
@@ 4.4 Explanations of $H_{0}$ and $S_{8}$ tensions
$H^{2}(z)=H_{0}^{2}\left[\Omega_{m}(1+z)^{3}+\Omega_{r}(1+z)^{4}+\Omega_{\text{TD}}(z)\right]$ (53)
@@
$\Omega_{\text{TD}}(z)=\frac{\rho_{\text{TD}}(z)}{3M_{P}^{2}H_{0}^{2}}$ (54)
@@
$D_{L}(z)=(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}$ (55)
@@
$\rho_{\text{TD}}(z)=\rho_{\text{TD},0}\left[1+\epsilon_{1}z+\epsilon_{2}z^{2}+\cdots\right]$ (56)
@@
$H^{2}(z)=H_{0}^{2}\left[1+\frac{3}{2}\Omega_{m}z+\Omega_{\text{TD},0}(\epsilon_{1}z+\epsilon_{2}z^{2}+\cdots)\right]$ (57)
@@
$\Psi_{\text{eff}}=\Psi+\alpha\varphi$
@@
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}[\bar{\rho}\delta+\rho_{\text{TD}}]$
@@
$\ddot{\delta}(k,a)+2H\dot{\delta}(k,a)-4\pi G\bar{\rho}[1+\Delta\mu(k,a)]\delta(k,a)=0$ (58)
@@
$S_{8}=\sigma_{8}\sqrt{\frac{\Omega_{m}}{0.3}}$ (59)
@@
$\bar{\Phi}(z)=\bar{\Phi}_{0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (60)
@@
$\rho_{\rm TD}(z)=\rho_{\rm TD,0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (61)
@@
$\frac{H^{2}(z)}{H_{0}^{2}}=\Omega_{m}(1+z)^{3}+\Omega_{r}(1+z)^{4}+\Omega_{\rm TD,0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (62)
@@
$(1+z)^{-n}\simeq 1-nz+\frac{n(n+1)}{2}z^{2}+\cdots$ (63)
@@
$\frac{H^{2}(z)}{H_{0}^{2}}\simeq 1+\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z+\mathcal{O}(z^{2})$ (64)
@@
$H(z)\simeq H_{0}\left[1+\frac{1}{2}\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z\right]$ (65)
@@
$\frac{1}{H(z)}\simeq\frac{1}{H_{0}}\left[1-\frac{1}{2}\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z\right]$ (66)
@@
$H_{0}^{\rm fit}\simeq H_{0}^{\rm true}\left[1+\frac{1}{4}\epsilon\,n\,\Omega_{\rm TD,0}\langle z\rangle\right]$ (67)
@@
$\Delta\mu(k,z)=-\mu_{0}\,(1+z)^{-p}\exp\left[-\left(\frac{k}{k_{*}}\right)^{2}\right]$ (68)
@@
$\ddot{\delta}(k,z)+2H\dot{\delta}(k,z)-4\pi G\,\bar{\rho}(z)\left[1+\Delta\mu(k,z)\right]\delta(k,z)=0$ (69)
@@
$s\simeq 1+\frac{3}{5}\Delta\mu_{\rm eff}$ (70)
@@
$\ln\left(\frac{D_{\rm TD}}{D_{\Lambda{\rm CDM}}}\right)\simeq\frac{3}{5}\Delta\mu_{\rm avg}\ln 2$ (71)
@@
$\Phi(z,\mathbf{x})=\bar{\Phi}_{0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]+\mu_{0}(1+z)^{-p}\int d^{3}k\,e^{-(k/k_{*})^{2}}\,\delta(k,z)\,e^{i\mathbf{k}\cdot\mathbf{x}}$ (73)

### v2

@@ 5 Cosmological Teleodynamics Framework
$\mathcal{P}[\Gamma]\propto\exp[-\beta\mathcal{A}[\Gamma]-\alpha\mathcal{K}[\Gamma]]$ (5.1)
@@
$\mathcal{A}[\Gamma]=\int dt\left[\frac{p^{2}}{2a^{2}m}+m\Psi(x,t)\right]$ (5.2)
@@
$\mathcal{K}[\Gamma]=\int dt\,\Phi\left(x(t),p(t);f,\mathcal{E}\right)$ (5.3)
@@
$\dot{x}=\frac{p}{a^{2}m}$ (5.4)
$\dot{p}=-\nabla\Psi(x,t)-\alpha\nabla\Phi(x,p,t)-Hp$ (5.5)
@@
This is helpful as it shows that the dynamics experiences an additional force proportional to the gradient of the bias functional. This force is not sourced from new matter fields but rather represents a coarse-grained structural response of the system to its own past evolution and environment. As a result, the presence of $\alpha\nabla\Phi$ alters the gravitational streaming of matter, modifies the effective potential, and influences the background expansion. By promoting the trajectory equations to a kinetic description, one obtains the teleodynamic Boltzmann equation, from which the modified continuity, Euler, and growth equations follow.
In a cosmological setting, the bias functional $\Phi(t,x)$ enters the maximum–caliber weight through the history action as an additional structural term influencing the statistical evolution of cosmic phase space. Since the universe on large scales is well described by an FLRW background with small fluctuations about homogeneity and isotropy, it is natural to separate the functional into its spatially averaged component and its inhomogeneous deviations, and so we can write
$\Phi(t,x)=\bar{\Phi}(t)+\varphi(t,x)$ (3.1)
It is important to understand how and why this decomposition works. It follows the same logic that underlies the treatment of any scalar field or effective quantity in cosmology, wherein the spatial average of the field governs the background expansion, while its fluctuations source clustering and local gravitational potentials. Because the teleodynamic bias functional can depend nonlocally on the matter distribution, the tidal environment, or the accumulated memory of the cosmic web, it can vary both temporally and spatially. One notes that the FLRW symmetries imply that only the spatial average of such a quantity can enter the background equations, while its space-dependent part must enter the perturbed sector. So, the decomposition (3.1) is not an assumption, but a direct consequence of applying the standard cosmological splitting to a quantity that affects both the background and perturbations.
Let us consider this split clearly, beginning with the Einstein–Hilbert action for gravity coupled to matter. We include the teleodynamic bias as an additive functional in the effective action obtained from coarse-graining over microscopic histories, and the result of this can be written schematically as
$S_{\rm eff}[g_{\mu\nu},\Psi_{m}]=\int d^{4}x\sqrt{-g}\left[\frac{M_{P}^{2}}{2}R+\mathcal{L}_{m}(g_{\mu\nu},\Psi_{m})\right]-\alpha\int d^{4}x\sqrt{-g}\,\Phi[g_{\mu\nu},\Psi_{m};\mathcal{E}]$ (5.6)
where $g_{\mu\nu}$ is the spacetime metric, $R$ is the Ricci scalar, $M_{P}$ is the reduced Planck mass, $\Psi_{m}$ collectively denotes the degrees of freedom of matter and $\mathcal{L}_{m}$ is the Lagrangian matter density, and the parameter $\alpha$ controls the strength of the teleodynamic contribution. The last term arises from the coarse-grained limit of the maximum–caliber path weight and represents the macroscopic imprint of nonlocal memory on the spacetime dynamics. $\Phi$ is the teleodynamic bias functional, which encodes a coarse-grained imprint of persistence, memory, and environment dependence, and the symbol $\mathcal{E}$ denotes collective environmental variables or invariants that can be constructed from the distribution of matter. We can now vary this action with respect to $g_{\mu\nu}$, which gives us modified Einstein equations
$M_{\rm P}^{2}G_{\mu\nu}=T_{\mu\nu}^{\rm m}+T_{\mu\nu}^{\rm TD}$ (5.7)
where the teleodynamic stress tensor is given as
$T_{\mu\nu}^{\rm TD}=-\frac{2}{\sqrt{-g}}\,\frac{\delta}{\delta g^{\mu\nu}}\Biggl(\alpha\int d^{4}x\sqrt{-g}\,\Phi(t,x)\Biggr)=\alpha\,\Phi\,g_{\mu\nu}-2\alpha\,\frac{\delta\Phi}{\delta g^{\mu\nu}}$ (5.8)
We note here that the first term in (5.8) behaves as an effective perfect-fluid contribution, while the second term encodes the possible metric dependence of the bias functional associated with tidal correlations or geometric memory. If $\Phi$ depends only on cosmic time at the background level, then $\delta\Phi/\delta g^{\mu\nu}=0$ for the homogeneous mode, and we obtain a perfect fluid form, which is
$T^{\mu}{}_{\nu\,{\rm TD}}={\rm diag}[-\rho_{\rm TD},\,p_{\rm TD},\,p_{\rm TD},\,p_{\rm TD}]$ (5.9)
with
$\rho_{\rm TD}(t)=\alpha\,\bar{\Phi}(t),\qquad p_{\rm TD}(t)=\alpha\,\bar{\Pi}_{\rm TD}(t)$ (5.10)
where $\bar{\Pi}_{\rm TD}(t)$ arises from the time dependence of $\bar{\Phi}(t)$ when inserted into the covariant conservation equation $\nabla_{\mu}T^{\mu\nu}_{\rm TD}=0$. These expressions show that the homogeneous component $\bar{\Phi}(t)$ behaves as an additional energy density and pressure at the background level. This can eventually modify the Friedmann equations while respecting the cosmological symmetries, as we shall see soon. This establishes clearly why a spatially averaged bias functional must influence the expansion history, as only its homogeneous mode survives in the background of an FLRW universe.
Now we would like to consider the inhomogeneous part $\varphi(t,x)$, which affects the perturbed sector and therefore determines the local clustering behavior. In the Newtonian gauge, the perturbed Einstein equations imply that the Poisson equation for the potential $\Psi$ becomes
$\nabla^{2}\Psi=4\pi Ga^{2}\bigl(\bar{\rho}\,\delta+\delta\rho_{\rm TD}\bigr)$ (5.11)
with
$\delta\rho_{\rm TD}(t,x)=\alpha\,\varphi(t,x)+\text{metric response terms}$ (5.12)
When the metric dependence of the bias functional is negligible on subhorizon scales, the dominant contribution can be identified as being
$\delta\rho_{\rm TD}(t,x)\simeq\alpha\,\varphi(t,x)$ (5.13)
This will act as an emergent clustering density, as we shall see later. In Fourier space, this becomes
$\delta\rho_{\rm TD}(k,t)=\frac{\alpha}{4\pi Ga^{2}}\,k^{2}\varphi_{k}(t)$ (5.14)
so that the effective potential satisfies
$\nabla^{2}\Psi_{\rm eff}=\nabla^{2}\bigl(\Psi+\alpha\,\varphi\bigr)=4\pi Ga^{2}\bigl(\bar{\rho}\,\delta+\delta\rho_{\rm TD}\bigr)$ (5.15)
This shows that the inhomogeneous component of the bias functional provides an additional source of gravitational potential, modifying galaxy motions, etc., thereby demonstrating that the homogeneous and inhomogeneous parts of $\Phi$ play distinct physical roles. While $\bar{\Phi}(t)$ governs the background expansion through its contribution to the stress tensor, $\varphi(t,x)$ shapes the spatial variations in the effective potential and thus modifies clustering. This division is unavoidable once one acknowledges that memory, environmental dependence, and persistent structure necessarily generate both large-scale coherent contributions and spatially varying corrections. The decomposition (3.1) is therefore a natural and physically required feature of any teleodynamic description of cosmology.Please note that extended derivations, clarifications, and a deeper discussion on the fundamental aspects of the teleodynamic action, the field decomposition, its perturbations, and gauge invariant properties are in the appendix. The appendix also clearly establishes how Cosmological Teleodynamics produces a cosmology which is fundamentally distinct from a simple barotropic fluid cosmology.
Also note that in the teleodynamic interpretation developed in the following sections, the "agents" of the cosmic game are not individual particles, stars, or local structures, but galaxies understood as coarse-grained units that propagate long-lived dynamical memory. This choice follows naturally from the fact that galaxies, rather than microscopic constituents, are the fundamental carriers of persistent organization in the late Universe as they preserve tidal histories, respond coherently to their large-scale environments, and act as the basic building blocks whose collective interactions generate the cosmic web. Treating galaxies as agents, therefore, reflects their role as the minimal scale at which non-ergodicity, correlation persistence, and teleodynamic bias manifest in a physically meaningful way. Thus, galaxies belong to the third category of agents, defined in statistical teleodynamics as intrinsically persistent systems whose behavior arises from physical constraints, such as long-range gravitational memory and accumulated structure, rather than from intention or cognition. Therefore, galaxies participate in a potential game-like evolution, with their aggregate dynamics shaping the emergent teleodynamic forces that drive various aspects of cosmology, as we shall see below.
@@ 5.1 Boltzmann Dynamics and Galactic Phase-space Evolution
$\frac{\partial f}{\partial t}+\frac{\mathbf{p}}{a^{2}m}\cdot\nabla_{\mathbf{x}}f-\nabla_{\mathbf{x}}\Psi\cdot\nabla_{\mathbf{p}}f-H\,\mathbf{p}\cdot\nabla_{\mathbf{p}}f=0$ (5.16)
@@
$\Phi(t,\mathbf{x})=\bar{\Phi}(t)+\varphi(t,\mathbf{x})$ (3.1)
@@
$\mathbf{a}=-\nabla\Psi-\alpha\,\nabla\Phi$ (5.17)
@@
$\frac{\partial f}{\partial t}+\frac{\mathbf{p}}{a^{2}m}\cdot\nabla_{\mathbf{x}}f-\left(\nabla_{\mathbf{x}}\Psi+\alpha\nabla_{\mathbf{x}}\Phi\right)\cdot\nabla_{\mathbf{p}}f-H\,\mathbf{p}\cdot\nabla_{\mathbf{p}}f=0$ (5.18)
@@
$\dot{\delta}+\frac{1}{a}\nabla\cdot\left[(1+\delta)u\right]=0$ (5.19)
@@
$\dot{u}+Hu+\frac{1}{a}(u\cdot\nabla)u=-\frac{1}{a}\nabla\Psi-\frac{1}{a(1+\delta)}\nabla\cdot P-\frac{\alpha}{a}\nabla\Phi$ (5.20)
@@
$\dot{\delta}(k,t)=-\frac{1}{a}\,\theta(k,t),$ (5.21)
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga\bar{\rho}\,\delta(k,t)=-\frac{c_{s}^{2}k^{2}}{a}\,\delta(k,t)-\frac{\alpha k^{2}}{a}\,\Phi_{k}(t)$ (5.22)
@@ 5.2 Friedmann Dynamics and Dark Energy
$\dot{\rho}+3H(\rho+p)=0,$ (5.26)
these relations fully determine the FLRW background given a set of matter components with equations of state. One immediate consequence of (5.24)-(5.25) is that cosmic acceleration
$\frac{\ddot{a}}{a}>0,$ (5.27)
@@
$\rho_{\text{TD}}(t)=\alpha\bar{\Phi}(t),\quad p_{\text{TD}}(t)=\alpha\Pi_{\text{TD}}(t)$ (5.30)
Note that no additional fields are required for this term to appear. It follows directly from the statistical and systemic coarse-graining of matter in a gravitationally correlated environment and represents the macroscopic imprint of dynamical memory and nonlocal structure. One can derive (5.28) and (5.29) directly by inserting the modified stress tensor
@@
$S_{\mathrm{eff}}[g_{\mu\nu}]=\int d^{4}x\,\sqrt{-g}\left[\mathcal{L}_{\mathrm{matter}}-\alpha\,\Phi\right]$ (5.31)
@@
$\frac{\ddot{a}}{a}=-\frac{1}{6M_{P}^{2}}[\rho+3p+\rho_{\text{TD}}+3p_{\text{TD}}]>0$ (5.32)
@@
$w_{\text{TD}}\equiv\frac{p_{\text{TD}}}{\rho_{\text{TD}}}=\frac{\Pi_{\text{TD}}(t)}{\bar{\Phi}(t)}\approx-1$ (5.33)
@@
Instead of introducing new scalar fields, potentials, or hidden matter sectors in our theory [53, 54, 55], one simply recognizes that the universe is not a system that relaxes to a gas-like equilibrium or one whose expansion is governed solely by local microphysics. The long-range nature of gravity and the persistent structures generated by the Big Bang and subsequent cosmic evolution imply that the background expansion should include contributions that have nonlocal correlations and path dependence. The teleodynamic energy density $\rho_{\text{TD}}$ arises precisely from these considerations and provides an emergent explanation for the cosmic acceleration. The term $\rho_{\mathrm{TD}}=\alpha\bar{\Phi}$ arises not from a new microscopic field, but from the coarse-grained effect of self-gravitating correlations and memory, which accumulate as the Universe forms persistent structures and departs more strongly from ergodicity. In this sense, we see that late-time acceleration emerges naturally as the Universe becomes increasingly non-thermal and the expansion is governed by an effective potential generated by the ensemble of long-lived gravitational correlations. Teleodynamics in this sense provides a statistical and systemic explanation for cosmic acceleration grounded in the physics of memory-bearing, nonlocally correlated matter rather than in additional dark sectors or fine-tuned vacuum energies.
@@ 5.3 Dark Matter and Structure Growth
$\Psi_{\text{eff}}(t,x)=\Psi(t,x)+\alpha\varphi(t,x)$ (5.34)
@@
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}\bar{\rho}\delta+\alpha\nabla^{2}\varphi$ (5.36)
@@
$\rho_{\text{TD}}(t,x)=\frac{\alpha}{4\pi Ga^{2}}\nabla^{2}\varphi(t,x)$ (5.37)
@@
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}[\bar{\rho}\delta+\rho_{\text{TD}}]$ (5.38)
@@
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga\bar{\rho}\delta(k,t)=-\frac{\alpha k^{2}}{a}\varphi_{k}(t)$ (5.39)
@@
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga[\bar{\rho}\delta(k,t)+\rho_{\text{TD}}(k,t)]=0$ (5.40)
@@ 5.4 Explanations of $H_{0}$ and $S_{8}$ tensions
$H^{2}(z)=H_{0}^{2}\left[\Omega_{m}(1+z)^{3}+\Omega_{r}(1+z)^{4}+\Omega_{\text{TD}}(z)\right]$ (5.43)
@@
$\Omega_{\text{TD}}(z)=\frac{\rho_{\text{TD}}(z)}{3M_{P}^{2}H_{0}^{2}}$ (5.44)
@@
$D_{L}(z)=(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}$ (5.45)
@@
$\rho_{\text{TD}}(z)=\rho_{\text{TD},0}\left[1+\epsilon_{1}z+\epsilon_{2}z^{2}+\cdots\right]$ (5.46)
@@
$H^{2}(z)=H_{0}^{2}\left[1+\frac{3}{2}\Omega_{m}z+\Omega_{\text{TD},0}(\epsilon_{1}z+\epsilon_{2}z^{2}+\cdots)\right]$ (5.47)
@@
$\Psi_{\text{eff}}(t,x)=\Psi(t,x)+\alpha\varphi(t,x)$ (5.34)
@@
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}[\bar{\rho}\delta+\rho_{\text{TD}}]$ (5.38)
@@
$\ddot{\delta}(k,a)+2H\dot{\delta}(k,a)-4\pi G\bar{\rho}[1+\Delta\mu(k,a)]\delta(k,a)=0$ (5.48)
@@
$S_{8}=\sigma_{8}\sqrt{\frac{\Omega_{m}}{0.3}}$ (5.49)
@@
$\bar{\Phi}(z)=\bar{\Phi}_{0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (5.50)
@@
$\rho_{\rm TD}(z)=\rho_{\rm TD,0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (5.51)
@@
$\frac{H^{2}(z)}{H_{0}^{2}}=\Omega_{m}(1+z)^{3}+\Omega_{r}(1+z)^{4}+\Omega_{\rm TD,0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (5.52)
@@
$(1+z)^{-n}\simeq 1-nz+\frac{n(n+1)}{2}z^{2}+\cdots$ (5.53)
@@
$\frac{H^{2}(z)}{H_{0}^{2}}\simeq 1+\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z+\mathcal{O}(z^{2})$ (5.54)
@@
$H(z)\simeq H_{0}\left[1+\frac{1}{2}\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z\right]$ (5.55)
@@
$\frac{1}{H(z)}\simeq\frac{1}{H_{0}}\left[1-\frac{1}{2}\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z\right]$ (5.56)
@@
$H_{0}^{\rm fit}\simeq H_{0}^{\rm true}\left[1+\frac{1}{4}\epsilon\,n\,\Omega_{\rm TD,0}\langle z\rangle\right]$ (5.57)
@@
$\Delta\mu(k,z)=-\mu_{0}\,(1+z)^{-p}\exp\left[-\left(\frac{k}{k_{*}}\right)^{2}\right]$ (5.58)
@@
$\ddot{\delta}(k,z)+2H\dot{\delta}(k,z)-4\pi G\,\bar{\rho}(z)\left[1+\Delta\mu(k,z)\right]\delta(k,z)=0$ (5.59)
@@
$s\simeq 1+\frac{3}{5}\Delta\mu_{\rm eff}$ (5.60)
@@
$\ln\left(\frac{D_{\rm TD}}{D_{\Lambda{\rm CDM}}}\right)\simeq\frac{3}{5}\Delta\mu_{\rm avg}\ln 2$ (5.61)
@@
$\Phi(z,\mathbf{x})=\bar{\Phi}_{0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]+\mu_{0}(1+z)^{-p}\int d^{3}k\,e^{-(k/k_{*})^{2}}\,\delta(k,z)\,e^{i\mathbf{k}\cdot\mathbf{x}}$ (5.63)